OpenStudy (readergirl12):

What is the solution to the system? Line 1 x y –1 –2 1 2 Line 2 x y –2 10 4 1

OpenStudy (studentchaitrali):

you wanna find equations?

OpenStudy (studentchaitrali):

Consider (-1,-2) as (x1,y1) and (1,2) as (x2,y2) and then substitute them in the eqn (x-x1)/(x1-x2)=(y-y1)/(y1-y2)

OpenStudy (readergirl12):

Okie.

OpenStudy (readergirl12):

(2,4)? @StudentChaitrali

OpenStudy (studentchaitrali):

You wanna find the straight line equation or wat?

OpenStudy (studentchaitrali):

The answer if you r finding an equation 2x-y=0